### v1

@@ 2 Medical Score Computing and its Ground Truth Labeling
Table 1: Mapping Formalism to a Concrete Example: The LACE Score
@@ 2.1 Mathematical Formalism
    a fallback undefined token denoted by $\mathsf{NA}$, indicating unextractability of a feature.
In sum, $x_{i}\in\mathcal{X}_{i}:=\mathbb{R}\cup\mathcal{C}\cup\{\mathsf{NA}\}$. Let $\bm{\mathcal{X}}\triangleq\prod_{i=1}^{m}\mathcal{X}_{i}$ denote the feature space, a Cartesian product of individual per-feature domains. Lastly, for each $(C,q)$ let $y$ be any answer predicted by an algorithm, and $y^{*}$ be the (latent) clinically correct ground truth answer; $y^{*}$ is a scalar if the question $q$ is answerable given $C$, and $\mathsf{NA}$ if $C$ lacks crucial information or is not clinically suited for answering $q$. Perfectly observing $y^{*}$ is costly, so an estimator $\hat{y}$ ($\hat{y}_{\textsf{original}}$ for MedCalc-Bench, $\hat{y}_{\textsf{new}}$ for ours) is used.
@@
where $\mathcal{Y}\subseteq\mathbb{R}\cup\{\mathsf{NA}\}$ is the space of possible score values. If at least one feature in $\bm{x}$ is out-of-range or $\mathsf{NA}$, the final score is not computable and $\mathsf{NA}$ is returned.
@@
Remark 2.1 (Insights from the Formalism)
The above formalism abstracts the sequential modularity of a typical medical computing workflow and the graphical structure of information flows: the first stage ($f_{\theta}$) interprets and extracts structured intermediate variables from unstructured data, while the second stage ($g_{\phi}$) performs a rule-based routine to aggregate the intermediate bits into a clinically verifiable and actionable scalar output. Of course, not all end-to-end workflows follow this specific pipeline structure; a system designer may as well eliminate $g_{\phi}$ and use a single call of $f_{\theta}$ to directly output the scalar output. That said, medcalc employed the former, structured data generation approach, so understanding this allows us to attribute any deviation in the final output $y$ to a parameter extraction error in $f_{\theta}$, or a formulaic mistake in $g_{\phi}$, or both.
2.2 Current task evaluation standard
To our best knowledge, the only large-scale public benchmark dataset for evaluating an agentic system’s end-to-end score computing capability is MedCalc-Bench (medcalc), created and released by researchers at NIH/NLM and five US medical schools. Their benchmark task asks an LLM model to read a patient note and answer a medical score computation question, and the model’s final answer $y$ is graded correct if it is within a $\pm 5\%$ margin of a gold label value $\hat{y}_{\textsf{original}}$ that the benchmark authors synthesized; and in cases where $y$ should be a datetime, exact string match is required. A model’s performance is quantified as the average accuracy on the benchmark’s test data set, which are 1047 instances manually reviewed by those authors.
@@ 2.2 Current task evaluation standard
    Patient notes are real, deidentified patient cases they scraped from journal-published medical papers that are archived in the PubMed Central database. There are altogether 10,053 + 1,047 = 11,100 notes in the train and test sets. We note that each of these patient notes can be seen as a unique context $C$ in the notation of §2.1.
@@
    $\hat{y}_{\textsf{original}}\in\mathcal{Y}$, MedCalc-Bench’s ground truth label of each $(C,q)$, was synthesized in a procedure analogous to the compositional pipeline we presented in 2.1. They used GPT-4 as $f_{\theta}$ and themselves hard-coded the logic of $g_{\phi}$ of each unique $q$ individually as 55 Python scripts.
@@
medcalc GPT-4 one-shot 50.9
oh2025 Llama3-70B 4-bit quantized 40.0
score-to-steps GPT-4o RAG + code 59.6*
riskagent GPT-4o multi-agent + tool use 67.7
med-u1 Qwen2.5-7B RLVR 57.6
ehrmind DeepSeek-R1 zero-shot 48.1
ehrmind Llama3-3B SFT + RLVR 52.0
@@
Since June 2024, multiple academic studies have used MedCalc-Bench as an evaluation benchmark or finetuned open-weight models on its train set. Table 2 shows a partial anthology of some test accuracies recently reported. Several survey papers on LLM and healthcare, including survey1; survey2; survey3, cited this benchmark. It is also included in the test suite of MedHELM (medhelm), a comprehensive framework for evaluating medical LLMs. These references indicate that MedCalc-Bench is an influential benchmark in the field.
2.3 Concerns about MedCalc’s Original Labeling Approach
MedCalc-Bench represents a significant contribution to evaluating AI for healthcare workflows. By sourcing real, de-identified patient notes and pairing them with 55 popular calculation tasks from MDCalc.com, the benchmark provides a far more realistic simulation of clinical tasks than traditional multiple-choice question answering. This curation represents a substantial effort by the authors to balance patient privacy with the need for accessible, high-quality evaluation data.
However, we have concerns about the procedure through which MedCalc-Bench synthesized the ground truth values $\hat{y}_{\textsf{original}}$’s, particularly regarding the reliability of their chosen $f_{\theta}$ and $g_{\phi}$. First, for both the train and test sets, the MedCalc authors used GPT4 as their feature extractor ($f_{\theta}$) to produce the clinical features vector $\bm{x}$, instructing it to output a JSON dictionary directly without intermediate reasoning. Although the authors reported manually checking this parameter extraction, their paper provides no statistics on the level of disagreement between their human reviewers and the GPT4-extracted values. Such disagreement seems probable, especially given that earlier versions of GPT4 were more prone to factual hallucinations than contemporary models.
Second, the level of alignment between their 55 Python functions (their $g_{\phi}$) and the formulas used by MDCalc.com is unclear. Given that MDCalc.com is used daily by many US physicians, its formulas represent a de facto standard. The authors did not detail the derivation process for their Python scripts, making it difficult to know if they were all fact-checked by licensed physicians.
As explained in Remark 2.1, each of these potential discrepancies in $f_{\theta}$ and $g_{\phi}$ could contribute to an accumulation of error in the final gold label. This motivates our work in §3 to systematically examine the true label quality of the benchmark.
Remark 2.2 (Aligning Ground Truth Labels)
@@ 2.3 Concerns about MedCalc’s Original Labeling Approach
3 Systematic Alignment of MedCalc Benchmark Labels
To address reliability concerns identified in §2.3 regarding the benchmark’s original labels, this section establishes a systematic framework for evaluating and improving label alignment with clinical knowledge. We organize our approach into two phases. In Phase 1 (§3.1), we conduct a scalable automated audit of the existing test set using a Gemini-2.5-Pro agentic pipeline, validated by a small-scale clinician review, to quantify the prevalence of label errors. Motivated by results uncovered in this audit, Phase 2 (§3.2) details a comprehensive relabeling workflow where we independently derive corrected ground truth values and then collect clinician feedback to quantify and compare the clinical alignedness of $\hat{y}_{\textsf{original}}$’s and $\hat{y}_{\textsf{new}}$’s.
Figure 1: In Phase 1, Gemini is endowed with Google search to individually validate 1047 $(C,q,\bm{y})$ instances provided in the original MedCalc-Bench release. After thinking and tool call(s), Gemini is instructed to give a binary judgment of if $\hat{y}_{\textsf{original}}$ is correct; the Gemini agent is queried 5 times, and we only count a case as problematic if there is supermajority agreement in the 5 responses. Note: $\bm{y}:=(\mathcal{M},\hat{y}_{\textsf{original}})$, where $\mathcal{M}$ is metadata (GPT4-extracted features $\bm{x}$ & text description of $g_{\phi}$) MedCalc-Bench provided for how $\hat{y}_{\textsf{original}}$ was derived.
3.1 Phase 1: Auditing Test Labels
Our pilot study focused on checking the label correctness of MedCalc’s test set, as this is the subset ultimately used to benchmark an LLM’s capability. The audit pipeline’s backbone uses the June 2025 version of Gemini-2.5-Pro (gemini25), which was one of the most advanced reasoning LLMs when we conducted the audit in July and August 2025. Figure 1 summarizes the structure of the audit pipeline. In words, Gemini is applied five times separately to clinically validate how the MedCalc label $\hat{y}_{\textsf{original}}$ was originally derived for each and every $(C,q)$, and we improve factual grounding by giving Gemini the ability to issue and reason on multiple Google search queries, rather than relying on its parametric knowledge. The reason for five applications of Gemini is to increase reliability, given the stochastic nature of LLMs generation. The overall pipeline design is partly motivated by the generation-verification gap (gv-gap); we believe that Gemini-2.5-Pro, enhanced with the said agentic abilities, serves as a strong verifier for outputs previously generated by a non-reasoning LLM in a single-turn, closed-book fashion (e.g. clinical features $\bm{x}$ extracted by GPT-4, or Python scripts codifying the $g_{\phi}$’s). Lastly, to reduce false positives, our system prompt asks Gemini to be conservative and only flag a test case as problematic if Gemini finds $\hat{y}_{\textsf{original}}$ to be incorrect in a clinically significant way. We provide the system prompt text and an example user prompt in the Appendix (Fig. 6).
3.1.1 Phase 1 Results
Figure 2: Phase 1 audit outcomes. 26.6% of MedCalc’s test instances are flagged as Likely Error’s by the Fig. 1 pipeline. A pilot clinician spot-check (7 flagged instances) agreed with the audit feedback in 7/7 instances, giving an early signal of the audit pipeline’s clinical validity.
Automated audit outcomes.
Running the pipeline on the full MedCalc test set (n=1,047n=1{,}047) surfaced a non-trivial number of label ($\hat{y}^{*}_{o}$) issues. In total, labels of 279/1,047 test instances (26.6%) were flagged as Likely Error by the verifier, while 768/1,047 (73.4%) passed (Fig. 2). Potential issues were distributed broadly across the distinct questions $q$ (55 calculators popular on MDCalc): in 40 of the 55, at least one associated test instance $(C,q)$ was flagged as a Likely Error. Since we only count an instance as Likely Error if $\geq 4/5$ independently sampled verifier runs all flag it, the statistics likely understate the true rate of controversial labels.
Physician expert feedback.
As a reality check on the Gemini auditor’s clinical judgment, a licensed physician independently selected and reviewed seven Likely Error instances related to his clinical specialty. For each instance, the physician read both $(C,q,\bm{y})$ and the Gemini auditor’s critique of $\bm{y}$. He agreed with the critiques in all 7/7 reviews. While the sample size is modest, this reality check supports the viability of using a reasoning LLM with tool use for label auditing.
Figure 3: Representative error types. (a) Feature extraction error, i.e. $f_{\theta}$ extracted a feature $x_{i}$ incorrectly: GPT-4 might have confused “hemoglobin” with “albumin”, scrapping a value that is physiologically impossible and biasing $\hat{y}_{\textsf{original}}$; (b) Incorrect aggregation logic ($g_{\phi}$): a Glasgow Coma Scale aggregation code that double-adds a feature value; (c) $q$ is not answerable given $C$ (mismatch / missing information): a Sodium correction for hyperglycemia applied to a clearly hypoglycemic patient. Row IDs for 3 examples in test.csv: 772, 849, 374.
Qualitative error taxonomy.
Manual inspection of flagged instances reveals three recurring failure modes in the original ground truth derivation: (a) feature extraction errors (flawed $f_{\theta}$), (b) incorrect aggregation logic (flawed $g_{\phi}$), and (c) $q$ is not answerable given $C$. Figure 3 illustrates one example of each. In these modes, (a) and (b) map directly onto the two-stage formalism in §2.1, whereas (c) correspond to either $f_{\theta}(C,q)=\mathsf{NA}$ or $g_{\phi}(\bm{x})=\mathsf{NA}$. In the case of (c), it would be more reasonable to reward the LLM for abstaining from answering the impossible and responding with $\mathsf{NA}$ (“Not Available”), instead of hallucinating an answer. Moreover, instead of throwing away all these instances, we suggest that the benchmark recycle these them by keeping $(C,q)$ but revising the ground truth label from $\hat{y}_{\textsf{original}}$ to $\mathsf{NA}$. Correspondingly, the benchmark’s inference prompt could acknowledge that there are instances where the correct answer is to abstain. A safely useful LLM system, we believe, should act as a triage agent that seeks human review for out-of-distribution $(C,q)$ instances, while strictly limiting abstention on normal cases to preserve the efficiency gains of automation.
3.2 Phase 2: Relabeling and Clinical Validation
Motivated by the findings in Phase 1, we proceeded to systematically relabel the dataset to establish a more reliable ground truth. We subdivided the Phase 2 experiment process into three parts: pipeline setup, automated error definition, and independent physician validation.
Pipeline 2 Setup.
We developed a separate relabeling pipeline (Fig. 4; for prompts we used, see Appendix Fig. 7) using a Gemini-2.5-Pro agent with access to Google Search and Python execution capabilities. Unlike Phase 1, the agent was provided only the patient context and question $(C,q)$, without access to the original MedCalc-Bench metadata including $\hat{y}_{\textsf{original}}$. We applied this pipeline to the full MedCalc test set (n=1,047n=1{,}047) and a stratified random sample of the training set (n=5,183n=5{,}183, 50% of total), stratified by question type $q$. To ensure label stability, each instance was processed by five independent copies this Gemini agent. If at least 4 of the 5 outputs agree in type and value (up to two decimal places for reals, exactly for datetimes, or all abstain), we keep that supermajority value as the new label $\hat{y}_{\textsf{new}}$; otherwise the instance is left unlabeled by Pipeline 2. This yields a set of “high-confidence” new labels.
Figure 4: Phase 2 labeling pipeline. The Gemini agent sees only $(C,q)$, has access to Google search and a code interpreter, and is asked to compute a label from first principles. For each instance we obtain five independent outputs $\hat{y}^{(1)}_{\textsf{new}},\dots,\hat{y}^{(5)}_{\textsf{new}}$. If at least four runs agree in type and value (up to two decimal places for reals, exact for datetimes or $\mathsf{NA}$), we keep the supermajority value as the new label $\hat{y}_{\textsf{new}}$; otherwise, the instance is deferred.
Flagging likely label errors.
Once a high-confidence $\hat{y}_{\textsf{new}}$ is collected from Pipeline 2 for a test instance $(C,q)$, we compare it to $\hat{y}_{\textsf{original}}$. We define a symmetric discrepancy metric
$\operatorname{rel\_err}(\hat{y}_{\textsf{original}},\hat{y}_{\textsf{new}}):=\frac{\left|\hat{y}_{\textsf{original}}-\hat{y}_{\textsf{new}}\right|}{\max\!\bigl(\left|\hat{y}_{\textsf{original}}\right|,\left|\hat{y}_{\textsf{new}}\right|\bigr)}.$
Because the denominator uses the maximum of the two estimates, $\operatorname{rel\_err}$ is a lower bound on the true relative error that would be obtained if either $\hat{y}_{\textsf{original}}$ or $\hat{y}_{\textsf{new}}$ exactly equaled the latent clinical ground truth $y^{\star}$. We classify an instance as Likely Error if $\operatorname{rel\_err}(\hat{y}_{\textsf{original}},\hat{y}_{\textsf{new}})>0.05$, mirroring the $\pm 5\%$ tolerance used by MedCalc-Bench for grading model outputs. Instances where either label is $\mathsf{NA}$ or datetime are handled respectively by exact match and >1>1 day difference.
Physician validation protocol.
To obtain an external clinical reference, we asked three licensed physicians (pediatrics, anesthesiology, oncology) to independently recompute a small subset of labels. The validation proceeds in three steps. First, we sort the potentially erroneous instances flagged by the above automated procedure in descending order of relative disagreement, counting $\mathsf{NA}$’s as the highest disagreement. Second, physicians screened the score questions $q$ to identify those matching their clinical specialty. Third, we selected 50 instances from the sorted list, assigning cases to reviewers based on domain alignment to ensure that every instance was validated by an expert in that specific calculator type. For each $(C,q)$ they review, a physician saw only $(C,q)$ and were asked to compute the score manually, without seeing either $\hat{y}_{\textsf{original}}$ or $\hat{y}_{\textsf{new}}$ (a single-blind design); physicians were allowed to label an instance as $\mathsf{NA}$ if they considered $(C,q)$ clinically unanswerable / a mismatch. Each of the 50 instances was labeled once by one of the three physicians, which we treat as the gold-standard answer $y^{\star}$. We then quantify the relative deviation of $\hat{y}_{\textsf{original}}$ and $\hat{y}_{\textsf{new}}$ against $y^{\star}$ by computing two metrics: (i) sMAPE on the instances a physician deemed answerable where sMAPE is short for symmetric Mean Absolute Percentage Error (armstrong1985), defined as
$\text{sMAPE}(y^{*},\hat{y})=\frac{100\%}{N}\sum_{i=1}^{N}\frac{2\times|y^{*}_{i}-\hat{y}_{i}|}{|y^{*}_{i}|+|\hat{y}_{i}|}\,$,
and (ii) count of “agreement” that counts, across all 50 instances, how often each synthetic label matches $y^{\star}$ within a clinically reasonable threshold, defined as $\pm 5\%$ for real numbers, exact match if $y^{*}=\mathsf{NA}$, or $\pm 1$ for ordinal numbers less than 20.
3.2.1 Phase 2 Results
Coverage and abstentions.
Pipeline 2 achieved high-confidence labels for 887 out of 1,047 MedCalc test instances (85%) and for 4,593 out of 5,183 sampled train instances (89%). Among these high-confidence labels, the abstention rate, cases where Pipeline 2 output $\mathsf{NA}$, was 7.4% on the test set and 5.7% on the sampled train set. Consistent with our discussion in §3.1, we retain such $\mathsf{NA}$ instances in the relabeled benchmark, as they correspond to clinically relevant “cannot be computed” cases rather than noise.
Estimated rate of likely mislabels.
Among the 887 test instances with high-confidence $\hat{y}_{\textsf{new}}$, we find that 220 (24.8%) exceed the $\pm 5\%$ numerical discrepancy threshold and are flagged as Likely Error under $\operatorname{rel\_err}$, plus another 66 (7.4%) are flagged as $\mathsf{NA}$. Together, this implies a total of 32.2% of the original test instances as likely mislabeled. We note that this is still likely an underestimate, as $\operatorname{rel\_err}$ is a conservative, symmetric lower bound on the underlying relative error with respect to the unknown $y^{\star}$. Moreover, the 160 test instances for which Pipeline 2 could not reach a supermajority (and are therefore excluded from this calculation) are plausibly hard cases where both $\hat{y}_{\textsf{original}}$ and $\hat{y}_{\textsf{new}}$ may be less reliable, potentially pushing the true mislabel rate above 32.2%.
Table 3: Original MedCalc test labels ($\hat{y}_{\textsf{original}}$) and our revised labels ($\hat{y}_{\textsf{new}}$) against 50 physician-generated labels ($y^{\star}$). sMAPE, short for symmetric Mean Absolute Percentage Error, reports the average relative percentage deviation from physicians’ labels; it is computed on 34 / 33 instances where the $y^{\star}$ and $\hat{y}_{\textsf{original}}$ / $\hat{y}_{\textsf{new}}$ are both numeric values. “Agreement” counts how many predicted labels in a group match the physician label up to a reasonable threshold. Square brackets are bootstrap $95\%$ confidence intervals.
Metric $\hat{y}_{\text{original}}$ $\hat{y}_{\text{new}}$
sMAPE (%) 72.7 [47.9, 99.1] 20.1 [8.0, 35.9]
Agreement 10/50 (20.0%) 37/50 (74.0%)
Physician validation.
Table 3 quantifies the prediction accuracy of $\hat{y}_{\textsf{original}}$ and $\hat{y}_{\textsf{new}}$ on the set of test instances with physician-computed labels, which we have treated as $y^{\star}$. Across the 50 instances, $\hat{y}_{\textsf{new}}$ agrees with $y^{\star}$ 37 times, compared to $\hat{y}_{\textsf{original}}$’s 10 times. Beyond binary agreement, the new labels drive sMAPE down from 72.7% to 20.1%, meaning that $\hat{y}_{\text{new}}$ typically lies much closer to physicians’ numerical assessments in both absolute and relative terms. Although the sample size is modest, these results provide a targeted reality check that the relabeled $\hat{y}_{\textsf{new}}$’s are more aligned with expert judgment on a substantial fraction of the contentious cases revealed by Pipeline 2.
3.2.2 Phase 2 Qualitative Analysis
Single-blind physician recomputations suggest that a substantial share of label disagreement is driven not by arithmetic mistakes, but by task underspecification: many instances do not contain enough contextual qualifiers to define a unique clinical “ground truth” score. Two patterns recur. First, the benchmark rarely specifies a scoring timepoint. Many notes describe multi-visit trajectories over weeks or months, but the question does not clarify whether the score should be computed at presentation, peak severity, discharge, or another reference time (e.g. rows 138, 334, 341, 509, 569, 739, 761, 785). Second, scoring rules can vary across time and jurisdiction, so the “official” aggregation rule (our $g_{\phi}$) is not uniquely describable by a static benchmark label. For example, Morphine Milligram Equivalent (MME) conversions for methadone depend on which guideline is assumed: pre-2023 CDC guidance used a tiered rule, i.e. a multiplier of 12 for methadone doses >60>60 mg/day (cdc_opioids_2016), whereas the current CDC guideline adopts a flatter multiplier (4.7) while cautioning that conversion becomes unreliable at higher doses (dowell2022cdc); some state-level guidance still reflects the earlier schedule (maryland_mme; utah_mme). Under either ambiguity, multiple defensible answers exist, so a single numeric gold label does not cleanly measure an LLM’s generalizable clinical capability.
Beyond underspecification, physician review also surfaced outright $(C,q)$ mismatches that induce a distribution shift: the patient context falls outside the population for which the calculator was derived or validated. A clear example is row 328, where the question asks for maintenance fluids for a 1.1 kg cat, despite the underlying dosing heuristic being estimated on humans. In such cases, forcing computation produces a clinically meaningless quantity. These instances are better treated as abstentions (NA) or excluded from accuracy-based evaluation, and the benchmark prompt should explicitly allow safe non-response when $(C,q)$ is clinically incompatible.
Practically, these failure modes motivate adding metadata (timepoint + guideline/version) and expanding the label space to include abstention.
4 When Labels Become Rewards and a Yardstick: A Controlled Study of Label Choice in Model Alignment
§3 argued that MedCalc-Bench’s original labels are often misaligned with clinical reality and introduced a relabeled version with closer agreement to physicians. Because this benchmark is also used to train LLMs, label alignment is not only an evaluation concern: it directly shapes the behavior of next-generation models that are optimized to match these labels.
The primary goal of this section is to investigate whether the label discrepancies documented above have a meaningful impact on downstream model training, or whether such noise is effectively “washed out” by the strength of signal in the data. This is not a foregone conclusion: when training on thousands of examples, one might expect that moderate label noise would be absorbed and have little effect on final model accuracy. Conversely, if label choice does materially affect trained model performance, this would underscore the practical importance of benchmark maintenance efforts like ours.
To test this, we view model alignment as a reinforcement learning (RL) stage in which a base model is iteratively updated to maximize a reward defined with respect to benchmark labels. Changing from MedCalc’s original labels $\hat{y}_{\textsf{original}}$ to our corrected labels $\hat{y}_{\textsf{new}}$ is therefore equivalent to changing the reward function and, by extension, the model’s implicit alignment objective. We run a controlled comparison study in which the same RL training pipeline is applied to Qwen3-8B with only the reward labels varied, holding all other factors constant. Crucially, in our experimental design, both the training rewards and the evaluation metric are defined with respect to the corrected labels $\hat{y}_{\textsf{new}}$. This means the comparison is not strictly apples-to-apples if one believes the two label sets differ meaningfully. However, this is precisely the point: if $\hat{y}_{\textsf{original}}$ and $\hat{y}_{\textsf{new}}$ were interchangeable up to inconsequential noise, we would expect both training conditions to yield similar test accuracy. A substantial performance gap would rule out this null hypothesis and demonstrate that the label discrepancies have real downstream consequences.
We emphasize that our goal is not to claim state-of-the-art benchmark performance. Rather, we seek to establish that label quality has tangible consequences for model alignment. That said, if one accepts the evidence from §3 that the corrected labels more closely reflect clinical ground truth, then a model trained on these labels would, in principle, be better aligned with physician judgment. We consider this a secondary implication and leave more rigorous validation of this claim—including broader physician evaluation of model outputs—to future work. The primary contribution of this section is simply to demonstrate that the choice of benchmark labels is not merely a measurement detail, but a consequential design decision that shapes the models we build.
4.1 Model alignment as policy optimization

### v2

@@ Appendix D Computing Medical Scores: a Formal Lens
Table 2: Mapping Formalism to a Concrete Example: The LACE Score[42]
@@ D.1 Mathematical formalism
    a fallback undefined token denoted by $\mathsf{N/A}$, indicating unextractability of a feature.
In sum, $x_{i}\in\mathcal{X}_{i}:=\mathbb{R}\cup\mathcal{C}\cup\{\mathsf{N/A}\}$. Let $\bm{\mathcal{X}}\triangleq\prod_{i=1}^{m}\mathcal{X}_{i}$ denote the feature space, a Cartesian product of individual per-feature domains. Lastly, for each $(C,q)$ let $y$ be any answer predicted by an algorithm, and $y^{*}$ be the (latent) clinically correct ground truth answer; $y^{*}$ is a scalar if the question $q$ is answerable given $C$, and $\mathsf{N/A}$ if $C$ lacks crucial information or is not clinically suited for answering $q$. Perfectly observing $y^{*}$ is costly, so an estimator $\hat{y}$ ($\hat{y}^{\mathrm{original}}$ for MedCalc-Bench, $\hat{y}^{\mathrm{new}}$ for ours) is used.
@@
where $\mathcal{Y}\subseteq\mathbb{R}\cup\{\mathsf{N/A}\}$ is the space of possible score values. If at least one feature in $\bm{x}$ is out-of-range or $\mathsf{N/A}$, the final score is not computable and $\mathsf{N/A}$ is returned.
@@
Insights from the Formalism
The above formalism abstracts the sequential modularity of a typical medical computing workflow and the graphical structure of information flows: the first stage ($f_{\theta}$) interprets and extracts structured intermediate variables from unstructured data, while the second stage ($g_{\phi}$) performs a rule-based routine to aggregate the intermediate bits into a clinically verifiable and actionable scalar output. Of course, not all end-to-end workflows follow this specific pipeline structure; a system designer may as well eliminate $g_{\phi}$ and use a single call of $f_{\theta}$ to directly output the scalar output. That said, [16] employed the former, structured data generation approach, so understanding this allows us to attribute any deviation in the final output $y$ to a parameter extraction error in $f_{\theta}$, or a formulaic mistake in $g_{\phi}$, or both.
D.2 Current task evaluation standard
To our best knowledge, the only large-scale public benchmark dataset for evaluating an agentic system’s end-to-end score computing capability is MedCalc-Bench, created and released by researchers at NIH/NLM and five US medical schools.[16] Their benchmark task asks an LLM model to read a patient note and answer a medical score computation question, and the model’s final answer $y$ is graded correct if it is within a $\pm 5\%$ margin of a gold label value $\hat{y}^{\mathrm{original}}$ that the benchmark authors synthesized; and in cases where $y$ should be a datetime, exact string match is required. A model’s performance is quantified as the average accuracy on the benchmark’s test data set, which are 1047 instances manually reviewed by those authors.
@@ D.2 Current task evaluation standard
    Patient notes are real, deidentified patient cases they scraped from journal-published medical papers that are archived in the PubMed Central database. There are altogether 10,053 + 1,047 = 11,100 notes in the train and test sets. We note that each of these patient notes can be seen as a unique context $C$ in the notation of §D.1.
@@
    $\hat{y}^{\mathrm{original}}\in\mathcal{Y}$, MedCalc-Bench’s ground truth label of each $(C,q)$, was synthesized in a procedure analogous to the compositional pipeline we presented in D.1. They used GPT-4 as $f_{\theta}$ and themselves hard-coded the logic of $g_{\phi}$ of each unique $q$ individually as 55 Python scripts.
@@
[16] GPT-4 one-shot 50.9
[28] Llama3-70B 4-bit quantized 40.0
[43] GPT-4o RAG + code 59.6*
[23] GPT-4o multi-agent + tool use 67.7
[47] Qwen2.5-7B RLVR 57.6
[21] DeepSeek-R1 zero-shot 48.1
[21] Llama3-3B SFT + RLVR 52.0
@@
Since June 2024, multiple academic studies have used MedCalc-Bench as an evaluation benchmark or finetuned open-weight models on its train set. Table 3 shows a partial anthology of some test accuracies recently reported. Several survey papers on LLM and healthcare, including [45, 48, 46], cited this benchmark. In addition, MedCalc-Bench is also included in the test suite of MedHELM,[6] a comprehensive framework for evaluating medical LLMs. These references indicate that MedCalc-Bench is an influential benchmark in the field. As all these studies were posted online before October 2025, i.e. before MedCalc-Bench authors had publicly recommended any changes to their dataset version in November 2025, it is fairly certain that their evaluations were based on the original NeurIPS version, i.e. (v1.0).
D.3 A statistical estimation lens of ground truth assurance
@@
D.4 Model alignment as policy optimization